# Prompt-Driven and Training-Free Forgetting Approach and Dataset for Large Language Models

### v1

@@ 1 Introduction
Existing research on model unlearning has focused on computer vision tasks like object removal and scene modification [9, 41, 39] and natural language processing [40, 36, 45], primarily relying on fine-tuning or pruning. For generative models, feature forgetting [12, 32] is crucial for precise removal of specific features while maintaining controllability. However, current methods struggle to balance forgetting accuracy and background preservation, often degrading generation consistency and affecting non-forgotten regions.
Need for a New Dataset. Most existing unlearning datasets rely on data poisoning techniques (e.g., BadNets [13, 27, 44, 35]) by injecting backdoors or noise into training data. However, these methods have limited applicability to generative and multimodal models due to their high-dimensional nature, where a single sample can influence the entire data distribution. Thus, a standardized dataset covering diverse modalities (real and synthetic images) and tasks is essential for robust unlearning evaluation.
@@ 2.1 Model Unlearning
Existing unlearning methods include fine-tuning, incremental updates, parameter pruning, and adversarial training, primarily applied in computer vision and natural language processing (NLP) tasks to remove specific data or features from models [40, 36, 46]. However, these methods typically require substantial parameter adjustment or retraining, especially for large-scale models, resulting in high computational costs and resource demands [2, 42]. Additionally, these methods often struggle to preserve generation quality while achieving target-specific unlearning, especially in high-dimensional generative models [23].
@@ 2.2 Diffusion Model Unlearning
Model Unlearning and Object Removal both aim to eliminate specific content from generated outputs, but their core objectives and methodologies differ significantly. Model Unlearning focuses on modifying the learned representations within a model to ensure that certain information is no longer accessible or influential during generation. This process typically involves techniques such as fine-tuning, gradient-based updates, or explicit data deletion [8, 14, 25]. It is primarily applied in privacy compliance scenarios, such as meeting GDPR requirements, ensuring that models do not retain specific training data or learned concepts. However, since model unlearning directly alters model parameters, it is challenging to precisely locate and remove specific data without affecting the overall performance of the model. The uncontrollability of forgetting remains a major challenge in this field.
In contrast, Object Removal operates at the instance level by utilizing mask-based inpainting methods to erase specific objects while preserving the surrounding background integrity [41, 3]. These methods typically rely on the image inpainting capabilities of diffusion models, such as Inst-Inpaint, for object removal. However, existing approaches still face challenges related to consistency and manual intervention, as inpainted regions in complex scenes often suffer from edge discontinuities and texture mismatches, affecting visual coherence, especially in high-resolution images [11]. Additionally, many object removal methods rely on manually provided masks or textual prompts, restricting automation and increasing user effort.
@@ 2.4 Challenges in Unlearning
Evaluation Metrics. Current evaluation metrics for generative models, such as Frechet Inception Distance (FID) [5, 24, 26], effectively measure the quality of generated images but fail to fully capture unlearning performance. Other CLIP-based metrics, including CLIP Distance and CLIP Maximum Mean Discrepancy (CMMD) [18, 41], assess semantic alignment in generated images but are limited in evaluating inner-outer region consistency and specific target removal effectiveness. The high dimensionality and data diversity of diffusion models add complexity to establishing unified evaluation standards [33, 37].
Datasets. Few public datasets are designed for model unlearning evaluation, and traditional datasets like ImageNet [6] and COCO [22] lack task diversity and complexity, making them inadequate for assessing selective unlearning in diffusion models [10, 17]. Gandikota et al. [9] emphasized the need for high-quality, task-diverse synthetic datasets for evaluating concept forgetting and object removal. However, real-world datasets like RORD [29], while useful for object removal, lack category labels, limiting their suitability for prompt-based feature unlearning tasks that require well-defined target concepts.
@@ 4.4 ForgetMe Dataset
Bird $<$bird$>$ 11,788 11,486 97.44%
Cat $<$cat$>$ 10,000 9,555 95.55%
Dog $<$dog$>$ 20,258 18,408 90.87%
ImageNet $<$imagenet$>$ 5,000 4,377 87.54%

### v3

# ForgetMe: Evaluating Selective Forgetting in Generative Models

@@ 1 Introduction
Existing research on model unlearning has focused on computer vision tasks like object removal and scene modification [24, 40, 60, 47, 59] and natural language processing [75, 56, 41], primarily relying on fine-tuning or pruning. For generative models, feature forgetting [53, 73, 48, 49, 31] is crucial for precise removal of specific features while maintaining controllability. However, current methods struggle to balance forgetting accuracy and background preservation, often degrading generation consistency and affecting non-forgotten regions.
Need for a New Dataset. Most existing unlearning datasets rely on data poisoning techniques (e.g., BadNets [58, 70, 39, 71, 74, 52, 51, 50]) by injecting backdoors or noise into training data. However, these methods have limited applicability to generative and multimodal models due to their high-dimensional nature, where a single sample can influence the entire data distribution. Thus, a standardized dataset covering diverse modalities (real and synthetic images) and tasks is essential for robust unlearning evaluation.
@@ 2.1 Model Unlearning
Existing unlearning methods include fine-tuning, incremental updates, parameter pruning, and adversarial training, primarily applied in computer vision and natural language processing (NLP) tasks to remove specific data or features from models [57, 45, yu2025guideline]. However, these methods typically require substantial parameter adjustment or retraining, especially for large-scale models, resulting in high computational costs and resource demands [2, 62]. Additionally, these methods often struggle to preserve generation quality while achieving target-specific unlearning, especially in high-dimensional generative models [27].
@@ 2.2 Diffusion Model Unlearning
Model Unlearning and Object Removal both aim to eliminate specific content from generated outputs, but their core objectives and methodologies differ significantly. Model Unlearning focuses on modifying the learned representations within a model to ensure that certain information is no longer accessible or influential during generation. This process typically involves techniques such as fine-tuning, gradient-based updates, or explicit data deletion [8, 14, 29, 13, 43]. It is primarily applied in privacy compliance scenarios, such as meeting GDPR requirements, ensuring that models do not retain specific training data or learned concepts. However, since model unlearning directly alters model parameters, it is challenging to precisely locate and remove specific data without affecting the overall performance of the model. The uncontrollability of forgetting remains a major challenge in this field.
In contrast, Object Removal operates at the instance level by utilizing mask-based inpainting methods to erase specific objects while preserving the surrounding background integrity [61, 3, 12, 46, 42]. These methods typically rely on the image inpainting capabilities of diffusion models, such as Inst-Inpaint, for object removal. However, existing approaches still face challenges related to consistency and manual intervention, as inpainted regions in complex scenes often suffer from edge discontinuities and texture mismatches, affecting visual coherence, especially in high-resolution images [11, 67, 65]. Additionally, many object removal methods rely on manually provided masks or textual prompts, restricting automation and increasing user effort.
@@ 2.4 Challenges in Unlearning
Evaluation Metrics. Current evaluation metrics for generative models, such as Frechet Inception Distance (FID) [5, 28, 30, 66], effectively measure the quality of generated images but fail to fully capture unlearning performance. Other CLIP-based metrics, including CLIP Distance and CLIP Maximum Mean Discrepancy (CMMD) [18, 61, 64], assess semantic alignment in generated images but are limited in evaluating inner-outer region consistency and specific target removal effectiveness. The high dimensionality and data diversity of diffusion models add complexity to establishing unified evaluation standards [37, 54, 34].
Datasets. Few public datasets are designed for model unlearning evaluation, and traditional datasets like ImageNet [6, 22] and COCO [26, 72, 44] lack task diversity and complexity, making them inadequate for assessing selective unlearning in diffusion models [10, 17, 25, 23]. Gandikota et al. [9] emphasized the need for high-quality, task-diverse synthetic datasets for evaluating concept forgetting and object removal. However, real-world datasets like RORD [33], while useful for object removal, lack category labels, limiting their suitability for prompt-based feature unlearning tasks that require well-defined target concepts.
@@ 4.4 ForgetMe Dataset
Bird <<bird>> 11,788 11,486 97.44%
Cat <<cat>> 10,000 9,555 95.55%
Dog <<dog>> 20,258 18,408 90.87%
ImageNet <<imagenet>> 5,000 4,377 87.54%